[Brahms-l] Shift report 20040302 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Tue Mar 02 2004 - 20:15:51 EST
------------------------------------------------------------
Date: Tue Mar  2 20:15:51 2004, Shift: 16:00-24:00
Supervisor: kh, with: 
Log book#15 pages 72 - 72

Shift Summary:
------------------------------------------------------------
Had a nice store until 17:20 at which point it was dumped to try to get high rates until beam studies would start at 20:00.  The other experiments apparently got more than they would have with keeping the other beam, but we "lost the gamble" because the RHIC clock appears to have been off.  See e-log for more details.

Runs achieved:
run    tr2    tr3
10731  40k   193k
10732  20k   100k

There were no other problems that I could see.

Things are shutdown now for beam studies.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Mar 2 20:16:05 2004

This archive was generated by hypermail 2.1.8 : Tue Mar 02 2004 - 20:16:27 EST